OpenStudy (gabbyalicorn):

Hitomi has a bag with four chips in it numbered 1 to 4. She randomly draws a chip and replaces it three times. What is the probability that Hitomi will draw a 1 all three times?

OpenStudy (misty1212):

HI! (again)

OpenStudy (gabbyalicorn):

Hello!

OpenStudy (misty1212):

what is the probability she draws a one on the first try?

OpenStudy (gabbyalicorn):

3/4 or 1/4 ?

OpenStudy (misty1212):

hmmm pick one (hint, there is only one "one")

OpenStudy (gabbyalicorn):

i think 1/4

OpenStudy (misty1212):

right!

OpenStudy (gabbyalicorn):

:D

OpenStudy (misty1212):

and since she replaces it, the probability she gets it on the second try is also \(\frac{1}{4}\)

OpenStudy (misty1212):

and also on the third try!

OpenStudy (gabbyalicorn):

so add 1/4 + 1/4 + 1/4?

OpenStudy (misty1212):

so all three, do \[\frac{1}{4}\times \frac{1}{4}\times \frac{1}{4}\]

OpenStudy (misty1212):

no dear don't ADD

OpenStudy (gabbyalicorn):

oh, multiply. Sorry.

OpenStudy (misty1212):

you want to multiply them together

OpenStudy (gabbyalicorn):

1/64

OpenStudy (misty1212):

it is an AND statement one and one and one

OpenStudy (misty1212):

right

OpenStudy (gabbyalicorn):

:D

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]

OpenStudy (gabbyalicorn):

Thank you!